OpenStudy (anonymous):

A volleyball is thrown up in the air with initial velocity of 8.2m/s. How long does it take the ball to reach maximum height? help please!

OpenStudy (zephyr141):

do you know the kinematic equations?

OpenStudy (zephyr141):

if you don't just google them.

OpenStudy (anonymous):

well i was given 3 to choose from. and i know im looking for the time. so im going to assume i have to use either, g=vf-vi/ t or y= vi(t)+1/2 (g) (t)^2

OpenStudy (zephyr141):

\[1:d=v_0t+\frac{1}{2}at^2\]\[2:v^2=v_0^2+2ad\]\[3:v=v_0+at\]\[4:d=\frac{v_0+v}{2}t\]

OpenStudy (zephyr141):

i would use number 3.

OpenStudy (anonymous):

oh then idk what i just gave you because thats like completely different but ok so how would i get t by itself

OpenStudy (zephyr141):

\[0=8.2+(-9.8)t\]\[-8.2=(-9.8)t\]\[\frac{-8.2}{-9.8}=t\]\[0.84s=t\]

OpenStudy (zephyr141):

subtract 8.2 from both sides. then divide both sides by (-9.8) to isolate t.

OpenStudy (zephyr141):

the first equation you gave me looks a lot like equation 4 that i gave you.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so then how do i know what the total time was that the ball was in the air?

OpenStudy (zephyr141):

well we found the time it took to get up to it's max height. assuming that nothing changes in it's descent then it should be the same amount of time on the way back down.